If you multiply Jon's age by 3 then take off 1 it is equal to jon's age plus 13 Find
his age

Answer by rfer(16322)

x+13=3x-1
x-3x=-1-13
-2x=-14
x=-14/-2
x=7 age
-----------------
7+13=3*7-1
20=20